A population consists of five members. The marital status of each member is given below, where M and S stand for married and single respectively.

 

Member

1

2

3

4

5

Marital Status

M

S

M

S

S


1
Expert's answer
2022-01-14T10:42:07-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \text{Members} & 1 & 2 & 3 & 4 & 5 \\ \hline
 \text{Marital status} & M & S & M & S & S \\
 \hdashline
\end{array}"

Find the proportion "P" of married members in population.


"P=\dfrac{2}{5}"

Select all possible sample of 2 members from the population without replacement.


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \text{Members} & 1 & 2 \\ \hline
 \text{Marital status} & M & S \\
 \hdashline
\end{array}, P_{12}=\dfrac{1}{2}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \text{Members} & 1 & 3 \\ \hline
 \text{Marital status} & M & M \\
 \hdashline
\end{array}, P_{13}=1"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \text{Members} & 1 & 4 \\ \hline
 \text{Marital status} & M & S \\
 \hdashline
\end{array}, P_{14}=\dfrac{1}{2}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \text{Members} & 1 & 5 \\ \hline
 \text{Marital status} & M & S \\
 \hdashline
\end{array}, P_{15}=\dfrac{1}{2}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \text{Members} & 2 & 3 \\ \hline
 \text{Marital status} & S & M \\
 \hdashline
\end{array}, P_{23}=\dfrac{1}{2}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \text{Members} & 2 & 4 \\ \hline
 \text{Marital status} & S & S \\
 \hdashline
\end{array}, P_{24}=0"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \text{Members} & 2 & 5 \\ \hline
 \text{Marital status} & S & S \\
 \hdashline
\end{array}, P_{25}=0"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \text{Members} & 3 & 4 \\ \hline
 \text{Marital status} & M & S \\
 \hdashline
\end{array}, P_{34}=\dfrac{1}{2}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \text{Members} & 3 & 5 \\ \hline
 \text{Marital status} & M & S \\
 \hdashline
\end{array}, P_{35}=\dfrac{1}{2}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \text{Members} & 4 & 5 \\ \hline
 \text{Marital status} & S & S \\
 \hdashline
\end{array}, P_{45}=0"

Compute the mean "\mu_P" of a sample proportion


"\mu_P=\dfrac{1}{10}(\dfrac{1}{2}+1+\dfrac{1}{2}+\dfrac{1}{2}+\dfrac{1}{2}+0+0"

"+\dfrac{1}{2}+\dfrac{1}{2}+0)=\dfrac{2}{5}"

Check

"\mu_P=\dfrac{2}{5}=P"

Compute the standard deviation "\sigma_P" of a sample proportion

"\sigma^2_P=\dfrac{1}{10}((\dfrac{1}{2}-\dfrac{2}{5})^2+(1-\dfrac{2}{5})^2"

"+(\dfrac{1}{2}-\dfrac{2}{5})^2+(\dfrac{1}{2}-\dfrac{2}{5})^2+(\dfrac{1}{2}-\dfrac{2}{5})^2"

"+(0-\dfrac{2}{5})^2+(0-\dfrac{2}{5})^2+(\dfrac{1}{2}-\dfrac{2}{5})^2"

"+(\dfrac{1}{2}-\dfrac{2}{5})^2+(0-\dfrac{2}{5})^2)=\dfrac{9}{100}"

"\sigma_P=\sqrt{\sigma^2}=\sqrt{\dfrac{9}{100}}=\dfrac{3}{10}"

Check


"\sqrt{\dfrac{P(1-P)}{n}\cdot\dfrac{N-n}{N-1}}=\sqrt{\dfrac{\dfrac{2}{5}(1-\dfrac{2}{5})}{2}\cdot\dfrac{5-2}{5-1}}"

"=\dfrac{3}{10}=\sigma_P"

"\sigma_P=\dfrac{3}{10}=\sqrt{\dfrac{P(1-P)}{n}\cdot\dfrac{N-n}{N-1}}"